OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (amistre64):

i need cramer practice :) x − 2y = 2 3x + y = 6 x = 6(-2)-2(1) / 3(-2) - 1(1) = -12 -2 / -6 -1 = 14/7 = 2 if i did it right

OpenStudy (amistre64):

by elimination we have: x − 2y = 2 6x +2y = 12 ------------ 7x = 14 x = 2